# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ Appendix C Proofs
Notation. Let $\mathcal{H}_{m}\subset\mathbb{R}_{+}^{m\times m}$ be a set of pairwise tree distance matrices for $m$ points. That is, given $\mathbf{H}\in\mathcal{H}_{m}$, the following properties satisfy: (i) there is a weighted tree $T=(V,E,\mathbf{A})$ such that $\mathbf{H}(j,j^{\prime})=d_{T}(j,j^{\prime})\;\forall\;j,j^{\prime}\in[m]$, (ii) $\mathbf{H}(j,j)=0\;\forall\;j\in[m]$, (iii) $\mathbf{H}(j_{1},j_{2})\leq\mathbf{H}(j_{1},j_{3})+\mathbf{H}(j_{3},j_{2})\;\forall\;j_{1},j_{2},j_{3}\in[m]$, and (v) $\mathbf{H}(j,j^{\prime})=\mathbf{H}(j^{\prime},j)\;\forall\;j,j^{\prime}\in[m]$. Let $\mathcal{W}_{m\leadsto n}\subset\mathbb{R}_{+}^{n\times n}$ represent the set of pairwise OT distance matrices for $n$ points, where the ground metric is given by the tree distance matrix $\mathbf{H}\in\mathcal{H}_{m}$ with $m$ points. That is, given $\mathbf{W}\in\mathcal{W}_{m\leadsto n}$, the following properties satisfy: (i) $\mathbf{W}(i,i^{\prime})=\mathrm{OT}(\bm{\mu}_{i},\bm{\mu}_{i^{\prime}},\mathbf{H})=\mathrm{TW}(\bm{\mu}_{i},\bm{\mu}_{i^{\prime}},T)$ for any two probability distributions $\bm{\mu}_{i},\bm{\mu}_{i^{\prime}}\in\mathbb{R}^{m}$ supported on the tree $T$, (ii) $\mathbf{W}(i,i)=0\;\forall\;i\in[n]$, (iii) $\mathbf{W}(i_{1},i_{2})\leq\mathbf{W}(i_{1},i_{3})+\mathbf{W}(i_{3},i_{2})$ for any probability distributions $\bm{\mu}_{i_{1}},\bm{\mu}_{i_{2}},\bm{\mu}_{i_{3}}\in\mathbb{R}^{m}$ supported on the tree $T$ , and (v) $\mathbf{W}(i,i^{\prime})=\mathbf{W}(i^{\prime},i)$ for any two probability distributions $\bm{\mu}_{i},\bm{\mu}_{i^{\prime}}\in\mathbb{R}^{m}$ supported on the tree $T$.
C.1 Proof of Theorem 4.1
Theorem 4.1.
@@
Consider the set $\mathcal{S}_{m}=\{\mathbf{H}_{c}\in\mathcal{H}_{m}|\left\lVert\mathbf{H}_{c}\right\rVert_{\infty}\leq\theta_{c}\text{ and }\mathbf{H}_{c}(j,j^{\prime})\geq\widetilde{\theta}_{c}\;\forall j\neq j^{\prime}\}$, the Wasserstein distance between samples using a hierarchical ground metric $\mathbf{H}_{c}$ is bounded by [Huizing et al., 2022]
@@
On this $\varepsilon$-subset of $\mathcal{K}$, the decoded binary tree is locally constant and thus continuous. Therefore, $F$ is continuous on $\mathcal{K}$ and $F$ is non-expansive [Bauschke et al., 2017]. Non-expansiveness, combined with bounded iterates, implies asymptotic regularity. By Opial’s lemma, every cluster point is a fixed point. Since the sequence is confined to the compact set $\mathcal{K}$, at least one limit point exists. Therefore, every limit point of the iteration is a fixed point, and the sequences generated by Alg. 1 admit at least one convergent subsequence whose limit is a fixed point of the alternating scheme.
@@
Since the resulting graph is fully connected and the node degrees are uniformly bounded below (i.e., $\deg(i)\geq d_{\min}$), the continuous-time diffusion operator $\mathbf{P}^{t}=\exp(-t\mathbf{L})$ satisfies the following heat kernel bound [Grigoryan, 2009]:
@@
For the lower bound, we use the inequality in Lemma C.1 and based on the result from [Leeb and Coifman, 2016, Lemma 3]:
@@
C.2 Proof of Theorem 4.2
Theorem 4.2.
@@
Because each of $\Psi({\mathbf{X}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)}))$ and $\Psi({\mathbf{Z}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)}))$ is firmly non-expansive, their composition is non-expansive. Based on [Bauschke et al., 2017], it yields asymptotic regularity. Together with boundedness, the asymptotic regularity activates Opial’s lemma, which guarantees that every cluster point of the sequence is a fixed point. The sequence $\widehat{\mathbf{W}}_{r}^{(l)}$ and $\widehat{\mathbf{W}}_{c}^{(l)}$ lies in the compact set. By Bolzano–Weierstrass it has at least one convergent sequence.
@@ Appendix D Incorporating learned hierarchical representation within HGCNs
We introduce an iterative learning scheme for joint hierarchical representation learning based on TWD, where the initial pairwise distance matrices $\mathbf{M}_{r}$ and $\mathbf{M}_{c}$ are typically computed using standard data-driven metrics, such as Euclidean distance or the distance based on cosine similarity. A key advantage of our approach is its ability to incorporate prior structural knowledge when a hierarchical structure is available for one of the modes. In such cases, the distance metric of the known hierarchy can be used to initialize the corresponding distance matrix, allowing the iterative scheme to refine the structure of the other mode. We demonstrate the applicability of this formulation to hierarchical graph data and show how it can be integrated into HGCNs [Chami et al., 2019, Dai et al., 2021].
In hierarchical graph data, the input is modeled as a hierarchical graph $H=([n],E,\mathbf{A})$, where $[n]=\{1,\ldots,n\}$ represents the nodes, and each node is associated with an $m$-dimensional attribute vector. We denote the collection of node features as a data matrix $\mathbf{X}\in\mathbb{R}^{n\times m}$, consistent with the high-dimensional data matrix setting in our framework. In hierarchical graph data, the node hierarchy (corresponding to the sample hierarchy in the data matrix) is known, while the feature structure remains implicit. Recent works have proposed developing graph convolutional networks in hyperbolic space [Chami et al., 2019, Liu et al., 2019, Dai et al., 2021, Zhang et al., 2021], where node features are projected from Euclidean to hyperbolic space.
@@ D.1 Hyperbolic feature transformation
It is important to note that our tree construction builds on the method proposed in [Lin et al., 2025], which decodes a binary tree from multi-scale hyperbolic embeddings (see App. A). As a result, after convergence, the produced hyperbolic embeddings can be directly used as initial hyperbolic node representations in HGCNs. For each node $i$, the multi-scale hyperbolic embedding is denoted by
$[(\mathbf{y}_{i}^{0})^{\top},\ldots,(\mathbf{y}_{i}^{K})^{\top}]^{\top},$ (44)
@@
The Poincaré half-space $\mathbb{H}^{m+1}$ is advantageous for representing the exponentially increasing scale of diffusion densities in hyperbolic space [Lin et al., 2023]. However, due to the computational advantages of performing Riemannian operations in the Lorentz model $\mathbb{L}^{m}$, and the known equivalence between the Poincaré and Lorentz models with diffeomorphism [Ratcliffe et al., 1994, Lin et al., 2024], we adopt the Lorentz model for feature transformation [Chami et al., 2019, Dai et al., 2021]. The hyperbolic feature transformation is performed by first transforming points from the Poincaré half-space to the Lorentz model, followed by applying the Lorentz linear transformation [Dai et al., 2021] and mapping back to the Poincaré half-space.
@@
The hyperbolic feature transformation is then applied via a Lorentz linear map [Dai et al., 2021]
@@ D.2 Hyperbolic neighborhood aggregation
where $[[i]]$ denotes the neighborhood of node $i$, $w_{is}$ are aggregation weights, and $\sigma$ is a non-linear activation function. In contrast to the Lorentz model formulation in [Chami et al., 2019, Dai et al., 2021], applying non-linear activations in the Poincaré half-space does not violate manifold constraints, as the operations remain in the same space. Notably, the weighted sum within the activation function corresponds to a weighted Riemannian mean in the Poincaré half-space model.
The integrated hyperbolic feature transformation and aggregation are similar to the formulation in [Dai et al., 2021], where all operations are conducted directly on the manifold. This contrasts with earlier methods such as [Chami et al., 2019], which rely on projecting points to the tangent space for linear operations and then mapping back. By operating entirely within the hyperbolic manifold, our integrated method maintains geometric fidelity throughout the learning process in HGCNs.
@@ Appendix E Details on experimental study
Dataset ##\## Samples / ##\## Nodes ##\## Classes ##\## Edges ##\## Features
@@ E.1 Datasets
Word-document benchmarks. We evaluate our method on four standard word-document benchmarks commonly used in the word mover’s distance [Kusner et al., 2015] and TWD literature [Yamada et al., 2022, Le et al., 2019, Chen et al., 2024, Lin et al., 2025]: (i) the BBCSPORT dataset, consisting of 13,243 bags of words (BOW) and 517 articles categorized into five sports types, (ii) the TWITTER dataset, comprising 6,344 BOW and 2,176 tweets classified into three types of sentiment, (iii) the CLASSIC dataset, including 24,277 BOW and 4,965 academic papers from four publishers, and (iv) the AMAZON dataset, containing 42,063 BOW and 5,600 reviews of four products. The pre-trained Word2Vec embeddings [Mikolov et al., 2013] are considered as the word embedding vector, trained on the Google News dataset, which includes approximately 3 million words and phrases. Word2Vec represents these words and phrases as vectors in $\mathbb{R}^{300}$. The document types serve as labels for classification tasks in Sec. 6.2.
Single-cell RNA-sequencing data. Two scRNA-seq datasets from [Dumitrascu et al., 2021] are considered:(i) the ZEISEL dataset: From the mouse cortex and hippocampus [Zeisel et al., 2015], comprising 4,000 gene markers and 3,005 single cells, and (ii) the CBMC dataset: from a cord blood mononuclear cell study [Stoeckius et al., 2017], consisting of 500 gene markers and 8,617 single cells. We used the divisive biclustering method in [Zeisel et al., 2015] to obtain 47 classes for Zeisel and 56 classes for the CBMC. The Gene2Vec [Du et al., 2019] is used as the gene embedding vectors [Bellazzi et al., 2021]. Cell types are used as classification labels in the classification experiments presented in Sec. 6.2.
Hierarchical graph datasets. For hierarchical graph data, we consider the following datasets: (i) the DISEASE dataset: constructed by simulating the SIR disease spreading model [Anderson and May, 1991], where node labels indicate infection status and node features indicate susceptibility to the disease [Chami et al., 2019], (ii) the AIRPORT dataset: a flight network dataset where nodes represent airports and edges represent airline routes. The label of a node indicates the population of the country where the airport is located [Chami et al., 2019], (iii) the CORA dataset: a citation network containing 2,708 nodes, 5,429 edges, and 1,433 features per node, with papers classified into seven machine learning categories [Sen et al., 2008], and (iv) the PUBMED dataset: Another citation network with 19,717 nodes, 44,338 edges, and 500 features per node, encompassing three classes of medicine publications [Sen et al., 2008]. The node labels are used for node classification experiments in Sec. 6.3.
@@ E.2 Hyperparameters settings
Initial distance metric. The initial pairwise distance matrices $\mathbf{M}_{r}\in\mathbb{R}^{n\times n}$ and $\mathbf{M}_{c}\in\mathbb{R}^{m\times m}$ are typically data-driven and constructed from the observations using standard similarity measures such as Euclidean or cosine distances [Lindenbaum et al., 2020]. These initial distances are then used to define a Gaussian kernel for the diffusion operator, as detailed in App.A. For word-document and scRNA-seq datasets, the initial distance matrices, $\mathbf{M}_{c}\in\mathbb{R}^{m\times m}$ (feature distances) and $\mathbf{M}_{r}\in\mathbb{R}^{n\times n}$ (sample distances), are computed using cosine similarity in the ambient space, following prior work [Jaskowiak et al., 2014, Kenter and De Rijke, 2015]. Empirically, cosine-based distances yield more stable and reliable tree representations compared to using Euclidean distances as initial distance, which were found to be less robust and led to decreased performance in our experiments. While different initial distance metrics lead to different tree representations, we observe that the resulting trees from Alg. 1 and Alg. 2 converge to unique representations when the regularization parameters $\gamma_{r}$ and $\gamma_{c}$ are sufficiently large (see App. F). However, we emphasize that such large regularization does not necessarily produce the most meaningful hierarchical representations. When $\gamma_{r}$ and $\gamma_{c}$ are too large, the regularization terms dominate the distance computation in the iterative learning scheme, leading to trees that may lack meaningful hierarchical structure.
@@
Hyperparameter configuration. We explore a broad range of hyperparameters to accommodate variations across tasks. For hyperparameter optimization, we conduct a grid search for each dataset using Optuna [Akiba et al., 2019]. The Gaussian kernel scale is selected from the set $\{0.1,1,2,5,10\}\times\chi$, where $\chi$ denotes the median of the pairwise distances [Zelnik-Manor and Perona, 2004, Ding and Wu, 2020]. The number of hyperbolic components in the product manifold is set with the range $K\in\{0,1,\ldots,19\}$. The regularization parameters $\gamma_{r}$ and $\gamma_{c}$ are chosen from the set $\{10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1},5\times 10^{-1},1,5,10^{1},5\times 10^{1},10^{2}\}$. The thresholds $\vartheta_{c}$ and $\vartheta_{r}$ in the filtering step are set as follows. At the first iteration ($l=0$), the total $L_{1}$ norms of the Haar coefficients across all samples and all features are computed, denoted by $\eta_{c}$ and $\eta_{r}$, respectively. Thresholds $\vartheta_{c}$ and $\vartheta_{r}$ are then respectively selected from the sets $\{0.1\eta_{c},0.2\eta_{c},\ldots,0.9\eta_{c}\}$ and $\{0.1\eta_{r},0.2\eta_{r},\ldots,0.9\eta_{r}\}$. As the optimal hyperparameter configuration varies across tasks and there is no universal choice that works for all cases, Optuna [Akiba et al., 2019] is used to systematically explore the parameter space and efficiently identify task-specific configurations.
Norm regularizer. We adopt a norm regularizer based on the snowflake penalty as in [Mishne et al., 2019], given by
[…]
@@
This smoothness-promoting regularizer increases monotonically over $[0,\infty)$ and imposes stronger penalties on small differences, thereby encouraging local regularity in the representations. While alternative regularization strategies, such as entropic regularization [Cuturi, 2013], could be considered, we demonstrate in App. F that they are less effective.

### v3

@@ Appendix C Proofs
Notation.
Let $\mathcal{H}_{m}\subset\mathbb{R}_{+}^{m\times m}$ be a set of pairwise tree distance matrices for $m$ points. That is, given $\mathbf{H}\in\mathcal{H}_{m}$, the following properties satisfy: (i) there is a weighted tree $T=(V,E,\mathbf{A})$ such that $\mathbf{H}(j,j^{\prime})=d_{T}(j,j^{\prime})\;\forall\;j,j^{\prime}\in[m]$, (ii) $\mathbf{H}(j,j)=0\;\forall\;j\in[m]$, (iii) $\mathbf{H}(j_{1},j_{2})\leq\mathbf{H}(j_{1},j_{3})+\mathbf{H}(j_{3},j_{2})\;\forall\;j_{1},j_{2},j_{3}\in[m]$, and (v) $\mathbf{H}(j,j^{\prime})=\mathbf{H}(j^{\prime},j)\;\forall\;j,j^{\prime}\in[m]$. Let $\mathcal{W}_{m\leadsto n}\subset\mathbb{R}_{+}^{n\times n}$ represent the set of pairwise OT distance matrices for $n$ points, where the ground metric is given by the tree distance matrix $\mathbf{H}\in\mathcal{H}_{m}$ with $m$ points. That is, given $\mathbf{W}\in\mathcal{W}_{m\leadsto n}$, the following properties satisfy: (i) $\mathbf{W}(i,i^{\prime})=\mathrm{OT}(\bm{\mu}_{i},\bm{\mu}_{i^{\prime}},\mathbf{H})=\mathrm{TW}(\bm{\mu}_{i},\bm{\mu}_{i^{\prime}},T)$ for any two probability distributions $\bm{\mu}_{i},\bm{\mu}_{i^{\prime}}\in\mathbb{R}^{m}$ supported on the tree $T$, (ii) $\mathbf{W}(i,i)=0\;\forall\;i\in[n]$, (iii) $\mathbf{W}(i_{1},i_{2})\leq\mathbf{W}(i_{1},i_{3})+\mathbf{W}(i_{3},i_{2})$ for any probability distributions $\bm{\mu}_{i_{1}},\bm{\mu}_{i_{2}},\bm{\mu}_{i_{3}}\in\mathbb{R}^{m}$ supported on the tree $T$ , and (v) $\mathbf{W}(i,i^{\prime})=\mathbf{W}(i^{\prime},i)$ for any two probability distributions $\bm{\mu}_{i},\bm{\mu}_{i^{\prime}}\in\mathbb{R}^{m}$ supported on the tree $T$.
C.1 Proof of Theorem 1
Theorem 1.
@@ C.1 Proof of Theorem 1
Consider the set $\mathcal{S}_{m}=\{\mathbf{H}_{c}\in\mathcal{H}_{m}|\left\lVert\mathbf{H}_{c}\right\rVert_{\infty}\leq\theta_{c}\text{ and }\mathbf{H}_{c}(j,j^{\prime})\geq\widetilde{\theta}_{c}\;\forall j\neq j^{\prime}\}$, the Wasserstein distance between samples using a hierarchical ground metric $\mathbf{H}_{c}$ is bounded by [59]
@@
On this $\varepsilon$-subset of $\mathcal{K}$, the decoded binary tree is locally constant and thus continuous. Therefore, $F$ is continuous on $\mathcal{K}$ and $F$ is non-expansive [140]. Non-expansiveness, combined with bounded iterates, implies asymptotic regularity. By Opial’s lemma, every cluster point is a fixed point. Since the sequence is confined to the compact set $\mathcal{K}$, at least one limit point exists. Therefore, every limit point of the iteration is a fixed point, and the sequences generated by Alg. 1 admit at least one convergent subsequence whose limit is a fixed point of the alternating scheme.
@@
Since the resulting graph is fully connected and the node degrees are uniformly bounded below (i.e., $\deg(i)\geq d_{\min}$), the continuous-time diffusion operator $\mathbf{P}^{t}=\exp(-t\mathbf{L})$ satisfies the following heat kernel bound [141]:
@@
For the lower bound, we use the inequality in Lemma C.1 and based on the result [55, Lemma 3]:
@@
C.2 Proof of Theorem 2
Theorem 2.
@@ C.2 Proof of Theorem 2
Because each of $\Psi({\mathbf{X}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)}))$ and $\Psi({\mathbf{Z}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)}))$ is firmly non-expansive, their composition is non-expansive, and it yields asymptotic regularity [140]. Together with boundedness, the asymptotic regularity activates Opial’s lemma, which guarantees that every cluster point of the sequence is a fixed point. The sequence $\widehat{\mathbf{W}}_{r}^{(l)}$ and $\widehat{\mathbf{W}}_{c}^{(l)}$ lies in the compact set. By Bolzano–Weierstrass it has at least one convergent sequence.
@@ Appendix D Incorporating learned hierarchical representation within HGCNs
We introduce an iterative learning scheme for joint hierarchical representation learning based on TWD, where the initial pairwise distance matrices $\mathbf{M}_{r}$ and $\mathbf{M}_{c}$ are typically computed using standard data-driven metrics, such as Euclidean distance or the distance based on cosine similarity. A key advantage of our approach is its ability to incorporate prior structural knowledge when a hierarchical structure is available for one of the modes. In such cases, the distance metric of the known hierarchy can be used to initialize the corresponding distance matrix, allowing the iterative scheme to refine the structure of the other mode. We demonstrate the applicability of this formulation to hierarchical graph data and show how it can be integrated into HGCNs [21, 45].
In hierarchical graph data, the input is modeled as a hierarchical graph $H=([n],E,\mathbf{A})$, where $n]=\{1,\ldots,n\}$ represents the nodes, and each node is associated with an $m$-dimensional attribute vector. We denote the collection of node features as a data matrix $\mathbf{X}\in\mathbb{R}^{n\times m}$, consistent with the high-dimensional data matrix setting in our framework. In hierarchical graph data, the node hierarchy (corresponding to the sample hierarchy in the data matrix) is known, while the feature structure remains implicit. Recent works have proposed developing graph convolutional networks in hyperbolic space [21, 94, 45, 142], where node features are projected from Euclidean to hyperbolic space.
@@ D.1 Hyperbolic feature transformation
It is important to note that the used tree construction [37] decodes a binary tree from multi-scale hyperbolic embeddings (see App. A). As a result, after convergence, the produced hyperbolic embeddings can be directly used as initial hyperbolic node representations in HGCNs. For each node $i$, the multi-scale hyperbolic embedding is denoted by
$(\mathbf{y}_{i}^{0})^{\top},\ldots,(\mathbf{y}_{i}^{K})^{\top}]^{\top},$ (44)
@@
The Poincaré half-space $\mathbb{H}^{m+1}$ is advantageous for representing the exponentially increasing scale of diffusion densities in hyperbolic space [50]. However, due to the computational advantages of performing Riemannian operations in the Lorentz model $\mathbb{L}^{m}$, and the known equivalence between the Poincaré and Lorentz models with diffeomorphism [110, 143], we adopt the Lorentz model for feature transformation [21, 45]. The hyperbolic feature transformation is performed by first transforming points from the Poincaré half-space to the Lorentz model, followed by applying the Lorentz linear transformation [45] and mapping back to the Poincaré half-space.
@@
The hyperbolic feature transformation is then applied via a Lorentz linear map [45]
@@ D.2 Hyperbolic neighborhood aggregation
where $[[i]]$ denotes the neighborhood of node $i$, $w_{is}$ are aggregation weights, and $\sigma$ is a non-linear activation function. In contrast to the Lorentz model formulation [21, 45], applying non-linear activations in the Poincaré half-space does not violate manifold constraints, as the operations remain in the same space. Notably, the weighted sum within the activation function corresponds to a weighted Riemannian mean in the Poincaré half-space model.
The integrated hyperbolic feature transformation and aggregation are conducted directly on the manifold [45]. This contrasts with the earlier method [21], which relies on projecting points to the tangent space for linear operations and then mapping back. By operating entirely within the hyperbolic manifold, our integrated method maintains geometric fidelity throughout the learning process in HGCNs.
@@ Appendix E Details on experimental study
Dataset #\# Samples / #\# Nodes #\# Classes #\# Edges #\# Features
@@ E.1 Datasets
Word-document benchmarks.
We evaluate our method on four standard word-document benchmarks commonly used in the word mover’s distance [25] and TWD literature [34, 35, 36, 37]: (i) the BBCSPORT dataset, consisting of 13,243 bags of words (BOW) and 517 articles categorized into five sports types, (ii) the TWITTER dataset, comprising 6,344 BOW and 2,176 tweets classified into three types of sentiment, (iii) the CLASSIC dataset, including 24,277 BOW and 4,965 academic papers from four publishers, and (iv) the AMAZON dataset, containing 42,063 BOW and 5,600 reviews of four products. The pre-trained Word2Vec embeddings [144] are considered as the word embedding vector, trained on the Google News dataset, which includes approximately 3 million words and phrases. Word2Vec represents these words and phrases as vectors in $\mathbb{R}^{300}$. The document types serve as labels for classification tasks in Sec. 6.2.
Single-cell RNA-sequencing data.
Two scRNA-seq datasets [84] are considered:(i) the ZEISEL dataset: From the mouse cortex and hippocampus [145], comprising 4,000 gene markers and 3,005 single cells, and (ii) the CBMC dataset: from a cord blood mononuclear cell study [146], consisting of 500 gene markers and 8,617 single cells. We used the divisive biclustering method [145] to obtain 47 classes for Zeisel and 56 classes for the CBMC. The Gene2Vec [147] is used as the gene embedding vectors [2]. Cell types are used as classification labels in the classification experiments presented in Sec. 6.2.
Hierarchical graph datasets.
For hierarchical graph data, we consider the following datasets: (i) the DISEASE dataset: constructed by simulating the SIR disease spreading model [78], where node labels indicate infection status and node features indicate susceptibility to the disease [21], (ii) the AIRPORT dataset: a flight network dataset where nodes represent airports and edges represent airline routes. The label of a node indicates the population of the country where the airport is located [21], (iii) the CORA dataset: a citation network containing 2,708 nodes, 5,429 edges, and 1,433 features per node, with papers classified into seven machine learning categories [9], and (iv) the PUBMED dataset: Another citation network with 19,717 nodes, 44,338 edges, and 500 features per node, encompassing three classes of medicine publications [9]. The node labels are used for node classification experiments in Sec. 6.3.
@@ E.2 Hyperparameters settings
The initial pairwise distance matrices $\mathbf{M}_{r}\in\mathbb{R}^{n\times n}$ and $\mathbf{M}_{c}\in\mathbb{R}^{m\times m}$ are typically data-driven and constructed from the observations using standard similarity measures such as Euclidean or cosine distances [148]. These initial distances are then used to define a Gaussian kernel for the diffusion operator, as detailed in App. A. For word-document and scRNA-seq datasets, the initial distance matrices, $\mathbf{M}_{c}\in\mathbb{R}^{m\times m}$ (feature distances) and $\mathbf{M}_{r}\in\mathbb{R}^{n\times n}$ (sample distances), are computed using cosine similarity in the ambient space, following prior work [149, 150]. Empirically, cosine-based distances yield more stable and reliable tree representations compared to using Euclidean distances as initial distance, which were found to be less robust and led to decreased performance in our experiments. While different initial distance metrics lead to different tree representations, we observe that the resulting trees from Alg. 1 and Alg. 2 converge to unique representations when the regularization parameters $\gamma_{r}$ and $\gamma_{c}$ are sufficiently large (see App. F). However, we emphasize that such large regularization does not necessarily produce the most meaningful hierarchical representations. When $\gamma_{r}$ and $\gamma_{c}$ are too large, the regularization terms dominate the distance computation in the iterative learning scheme, leading to trees that may lack meaningful hierarchical structure.
Hyperparameter configuration.
@@
We explore a broad range of hyperparameters to accommodate variations across tasks. For hyperparameter optimization, we conduct a grid search for each dataset using Optuna [151]. The Gaussian kernel scale is selected from the set $\{0.1,1,2,5,10\}\times\chi$, where $\chi$ denotes the median of the pairwise distances [152, 153]. The number of hyperbolic components in the product manifold is set with the range $K\in\{0,1,\ldots,19\}$. The regularization parameters $\gamma_{r}$ and $\gamma_{c}$ are chosen from the set $\{10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1},5\times 10^{-1},1,5,10^{1},5\times 10^{1},10^{2}\}$. The thresholds $\vartheta_{c}$ and $\vartheta_{r}$ in the filtering step are set as follows. At the first iteration ($l=0$), the total $L_{1}$ norms of the Haar coefficients across all samples and all features are computed, denoted by $\eta_{c}$ and $\eta_{r}$, respectively. Thresholds $\vartheta_{c}$ and $\vartheta_{r}$ are then respectively selected from the sets $\{0.1\eta_{c},0.2\eta_{c},\ldots,0.9\eta_{c}\}$ and $\{0.1\eta_{r},0.2\eta_{r},\ldots,0.9\eta_{r}\}$. As the optimal hyperparameter configuration varies across tasks and there is no universal choice that works for all cases, Optuna [151] is used to systematically explore the parameter space and efficiently identify task-specific configurations.
Norm regularizer.
We adopt a norm regularizer based on the snowflake penalty [57], given by
$\zeta(\mathbf{r}_{i}-\mathbf{r}_{i^{\prime}})=\frac{1}{2}\int_{0}^{\left\lVert\mathbf{r}_{i}-\mathbf{r}_{i^{\prime}}\right\rVert_{2}}\frac{1}{\sqrt{\xi}+10^{-6}}d\xi.$ (48)
@@
This smoothness-promoting regularizer increases monotonically over $0,\infty)$ and imposes stronger penalties on small differences, thereby encouraging local regularity in the representations. While alternative regularization strategies, such as entropic regularization [119], could be considered, we demonstrate in App. F that they are less effective.